OpenStudy (anonymous):

How long will it take $18,000.00 to grow to $21,000.00 if it is invested at 9%. Compounded quaterly??

hartnn (hartnn):

\(\large A =P (1+r/4)^{4n}\)

hartnn (hartnn):

A = amount after 'n' years P = principal (initial) amount n = number of years r = rate of interest

hartnn (hartnn):

can you find these values from the question?

hartnn (hartnn):

4n because quaterly compounded means interest is calculated 4 times a year

OpenStudy (anonymous):

ek minute. umm i can find the values of r, p & n

hartnn (hartnn):

you mean r,P, A ecause they are given in question you'll need to calculate n then

OpenStudy (anonymous):

what is nt

hartnn (hartnn):

in your formula n = number of times interest is calculated per year here, quaterly means n= 4 t is the time after which we get the final amount which needs to be calculated here

hartnn (hartnn):

i have just taken 'n' instead of 't' for number of years

hartnn (hartnn):

your formula a=p(1+r/n)^nt a =21000 p= 18000 r =9% = 0.09 n =4 find t

OpenStudy (anonymous):

ok i got it! thank youuu

hartnn (hartnn):

welcomeeee ^_^

OpenStudy (anonymous):

what did you get as t?

hartnn (hartnn):

1.731 lol, weird answer :P

OpenStudy (anonymous):

thats what i got too!!!! lol